OpenStudy (annie12m):

Is the set of numbers that can be written as the product of 6 and an integer closed under subtraction? A. Yes, because the difference of any two multiples of 6 is equal to 6. B. Yes, because the difference of any two multiples of 6 is also a multiple of 6. C. No, and a counterexample is 6 – 6 = 0. D. No, and a counterexample is 6 – 18 = –12.

OpenStudy (annie12m):

please help

OpenStudy (anonymous):

C and D are out right? since both 0 and -12 are multiples of 6

OpenStudy (jchick):

We need to get an understanding of what these number look like. >Is the set of numbers that can be written as the product of 6 and an integer closed The integers are {...-3, -2, -1. 0, 1, 2, 3, ...}

OpenStudy (jchick):

6 times each of those would be: {..., -18, -12, -6, 0, 6, 12, 18, ...}

OpenStudy (jchick):

any two of those numbers and subtract them, will you get an answer that is also in that set.

OpenStudy (jchick):

6-6 = 0

OpenStudy (jchick):

the difference of any two multiples of 6 is equal to 6. 12 is a multiple of 6 24 is a multiple of 6. 12 - 24 is not equal to 6. So, throw out option A. It is wrong.

OpenStudy (jchick):

6 - 6 = 0 which is a multiple of 6 because 0 times 6 = 0. So, throw out option C.

OpenStudy (jchick):

6 - 18 = -12 but -12 is a multiple of 6 because 6 times -2 = -12. Throw out option D

OpenStudy (jchick):

The answer is B because the difference of any two multiples of 6 is also a multiple of 6.

OpenStudy (annie12m):

so its ethier one of the yeses

OpenStudy (annie12m):

oh ok thanks

OpenStudy (jchick):

Well we eliminated A, C, and D

OpenStudy (jchick):

So that only left B

OpenStudy (jchick):

No problem!

OpenStudy (annie12m):

can you help me with 4 more questions?

OpenStudy (jchick):

Yes

OpenStudy (annie12m):

ok ill give you one at a time

OpenStudy (jchick):

Ok thanks!

OpenStudy (annie12m):

Numbers p and q are negative integers. Which statements are true? Choose all that are correct. A. p + q is a negative integer B. p • q is a positive integer C. p – q is a negative integer D. p/ q is a negative integer

OpenStudy (annie12m):

A is correct? B is wrong? C not sure. D is wrong?

OpenStudy (annie12m):

jckick you there?

OpenStudy (jchick):

let's assign p and q numbers p=3 and q=5

OpenStudy (jchick):

Yes

OpenStudy (jchick):

is 3+5 an integer? 3-5? 3*5? 3÷5?

OpenStudy (annie12m):

well it said p and q are both going to negative

OpenStudy (jchick):

A. p + q is a positive integer TRUE! Two positive numbers added = a positive number! Always! B. p – q is a negative integer Maybe not. Ie if P = 10 and q = 1, p-q = 10-1 = +9. FALSE C. p • q is a positive integer Yes. Two positive numbers multiplied together is still a positive integer D. is a positive integer

OpenStudy (annie12m):

so a and c are correct?

OpenStudy (jchick):

Well they said negative right?

OpenStudy (annie12m):

yeah

OpenStudy (annie12m):

so lets try putting it p and q negative in all the operations

OpenStudy (jchick):

Actually wait.

OpenStudy (jchick):

Just compare the two statements if they match then select it.

OpenStudy (annie12m):

what you mean?

OpenStudy (jchick):

c and d

OpenStudy (annie12m):

it goes in this order.. A. p + q is a negative integer B. p • q is a positive integer C. p – q is a negative integer D. p/ q is a negative integer

OpenStudy (jchick):

Well A is positive. so that is false

OpenStudy (annie12m):

you sure its not a and c?

OpenStudy (annie12m):

a negative plus a negative equals a negative

OpenStudy (annie12m):

Numbers p and q are negative integers.

OpenStudy (jchick):

Oops you are right I misread it. Sorry you are right the first one is negative.

OpenStudy (annie12m):

ok

OpenStudy (jchick):

Because if p = -2 + q = -2 you get - four.

OpenStudy (annie12m):

ya

OpenStudy (jchick):

Ok so how about multiplication?

OpenStudy (annie12m):

so is it a and c?

OpenStudy (jchick):

Tell me the rule for multiplying two negatives.

OpenStudy (annie12m):

multipulcation well a negative times a negative or a positive times a positve you get a positive and a negative time a positive or a positive times a negative you get a negative

OpenStudy (annie12m):

oh so b would be it too

OpenStudy (jchick):

two negatives make a positive

OpenStudy (jchick):

In multiplication.

OpenStudy (annie12m):

yeah

OpenStudy (jchick):

Ok I didn't understand what you had said.

OpenStudy (jchick):

So yes it would be B

OpenStudy (annie12m):

i said the same thing you said but i said it for all of them

OpenStudy (annie12m):

ok now we see subsraction

OpenStudy (jchick):

Wait

OpenStudy (jchick):

Subtracting a negative by a negative is a positive.

OpenStudy (jchick):

-2--2

OpenStudy (jchick):

This equals 0

OpenStudy (annie12m):

-6 - (-3) you keep the first number then switch it to addition then change the sign of the last number so now it is -6 plus 3 and that equals -3

OpenStudy (jchick):

Yes correct

OpenStudy (annie12m):

its still negative so its correct so c is also correct

OpenStudy (annie12m):

now lets try division

OpenStudy (jchick):

-8/-4 = +2

OpenStudy (annie12m):

thats wrong

OpenStudy (jchick):

If the 2 numbers you are multiplying or dividing have the SAME sign (+,+ or -,-) then the result is POSITIVE

OpenStudy (annie12m):

so its a, b, and c

OpenStudy (annie12m):

next question

OpenStudy (jchick):

Correct!

OpenStudy (annie12m):

Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (annie12m):

do you know what rational numbers are?

OpenStudy (jchick):

Ok so if you add any two elements of a set and you ALWAYS get another element in the set, that set is closed under addition. For example 2+2=4 2+4=6 and 4 and 6 are part of the even number set.

OpenStudy (annie12m):

so even would be correct

OpenStudy (jchick):

let's go through each set a) no 5 + 2 = 7. 7 is not in the set. b) yes c) yes d) 0+0=0

OpenStudy (annie12m):

so b c and d?

OpenStudy (annie12m):

but what are rational numbers?

OpenStudy (jchick):

B. even integers C. rational numbers D. {0}

OpenStudy (jchick):

a rational number is any number that can be expressed as the quotient or fraction p/q of two integers, p and q, with the denominator, q, not equal to zero.

OpenStudy (jchick):

In easy terms A number that can be made by dividing two integers.

OpenStudy (annie12m):

but what does that have to do with addition? so its not correct?

OpenStudy (jchick):

What I have posted is the correct answer.

OpenStudy (jchick):

Closure: The result of adding two rational numbers is another rational number.

OpenStudy (annie12m):

ok ill do those 3 as the answer then

OpenStudy (jchick):

Ok

OpenStudy (jchick):

Any more questions?

OpenStudy (annie12m):

now next question

OpenStudy (jchick):

Ok

OpenStudy (annie12m):

oh just 2 more left

OpenStudy (jchick):

Ok shoot.

OpenStudy (annie12m):

Which equations show that the set of whole numbers is not closed under subtraction? Choose all answers that are correct. A. 1 – (–2) = 3 B. 1 – 2 = –1 C. 2 – 0 = 2 D. 2 – 4 = –2

OpenStudy (annie12m):

oh and stay close attention to where it says NOT

OpenStudy (jchick):

Cave, in order to show closure for a set, you need to start with two elements in the set. (Keep in mind that negative numbers are NOT in the set of Whole numbers). So you need to look for which one starts with two whole numbers, then ends up with something that is not a whole number

OpenStudy (jchick):

It's not A

OpenStudy (annie12m):

b is correct?

OpenStudy (jchick):

option A shows us doing subtraction with a negative number. Negatives are not in the set of whole numbers. You need to start with this whole number - whole number =

OpenStudy (annie12m):

ok

OpenStudy (jchick):

Yes B is correct